# Deep Learning-based Intrusion Detection Systems: A Survey

### v1

@@ 5.2.3. Network Traffic-based Detectors
Network traffic comes from communications between hosts across a computer network. It is ruled by network protocols such as Transmission Control Protocol (TCP) and the User Datagram Protocol (UDP).
@@
ACO-SVM (Meng et al., 2020) is a network intrusion detection model that optimizes the parameters of SVM based on the ant colony algorithm. In this approach, the parameters of the SVM are regarded as the paths traversed by ants. Through collecting network status information, performing feature extraction and processing, and utilizing the pheromone update and node transfer of the ant colony algorithm to achieve path crawling, the optimal parameter combination of ACO-SVM is found. Min et al. (Min et al., 2021) propose an approach for network anomaly detection employing a memory-augmented deep autoencoder. This model learns the prototype patterns of normal inputs via the memory module and reconstructs the inputs of abnormal samples that are close to normal samples during training. Thus, even the latent vectors of abnormal inputs are transformed through the memory module and aggregated and reconstructed from similar normal samples learned from the memory that is learned only with normal samples, thereby resolving the issue of over-generalization. SigML++ (Trivedi et al., 2023) is a supervised anomaly detection approach featuring probabilistic polynomial approximation. It employs Fully Homomorphic Encryption and Artificial Neural Networks to approximate the sigmoid activation function, enabling the execution of anomaly detection without decrypting logs.
Recurrent Neural Network
HDLNIDS (Qazi et al., 2023) employs a Convolutional Recurrent Neural Network (CRNN) for the detection and classification of malicious network traffic, with the objective of enhancing the detection efficiency and accuracy of existing intrusion detection systems. It conducts local feature extraction via CNN and then utilizes the RNN layer to capture the temporal dependency of features, thereby enabling a more comprehensive identification of network traffic patterns.
5.2.4. Cross Log-based Detectors
Audit logs contain detailed system-level underlying information, network traffic logs depict interactions between systems, and application logs possess high-level semantic information. Employing various kinds of logs in DL-IDS can enhance the interpretability of logs and obtain a more comprehensive understanding of intrusions.
Traditional Learning
LogUAD (Wang et al., 2022b) is a method for offline detection. It does not require a log parsing procedure. LogUAD directly takes the original log message as input to avert the noise brought in by parsing. It uses Word2Vec to generate word vectors and combines Term Frequency-Inverse Document Frequency (TF-IDF) to generate weighted log sequence feature vectors for handling the evolution of log statements. Ultimately, it makes use of a computationally efficient unsupervised clustering algorithm to detect anomalies.
Recurrent Neural Network
LogAnomaly (Meng et al., 2019) constitutes a system log anomaly detection framework. The central notion of this framework is to model the log stream as a natural language sequence. LogAnomaly employs a pattern of offline learning and online detection. In the offline learning segment, LogAnomaly utilizes the FT-Tree to extract templates from historical logs and matches the historical logs with these templates. Subsequently, each template is transformed into a template vector sequence, and sequential and quantitative features are extracted via the LSTM model. In the online detection portion, LogAnomaly initially determines whether real-time logs can be matched with existing templates. If affirmative, it is converted into a template vector; otherwise, based on the similarity among template vectors, it will be approximated to an existing template vector. LogLS (Chen et al., 2022) is a system log anomaly detection approach based on symmetrically structured double-long short-term memory. LogLS is an optimization of the DeepLog method referred to in Section 5.2.2. It is designed to tackle the gradient issue of LSTM in long sequence prediction and accomplishes the prediction of unobserved logs by providing a feedback mechanism.
Transformer Architecture
Both NeuralLog (Le and Zhang, 2021) and LAnoBERT (Lee et al., 2023) are free from the need for a log parser. They are able to learn directly from the raw log data, thereby avoiding the issue of potentially losing crucial information during the standardization process. They both make use of BERT to encode the semantic significance of log messages and are capable of conducting online detection. The work presented in the paper LogDAPT (Zheng et al., 2023) is predominantly founded on NeuralLog. It is noted that NeuralLog necessitates a substantial amount of labeled data to yield satisfactory outcomes. In contrast, LogDAPT tackles the issue of insufficient labeled data by applying the second stage of in-domain pre-training, namely DAPT, on the BERT base model. Not only LogBERT (Guo et al., 2021) but also LogBASA (Liao et al., 2023) employ the log parser Drain to parse log events. Subsequently, they make use of the pre-trained BERT model to extract the semantic information of the events. In the TCN-Log2Vec (Jiang et al., 2023) framework, the log parsing component is optimized on the basis of Drain. Meanwhile, the BERT model is employed to precisely extract semantics from log templates. Additionally, TF-IDF is utilized to capture the context of log templates, and window-based techniques are adopted to obtain log sequence information. The ASGNet (Yang et al., 2023a) model is composed of three components: log statistics information representation (V-Net), log deep semantic representation (S-Net), and an adaptive semantic threshold mechanism (G-Net). V-Net employs an unsupervised variational autoencoder (VAE) to learn the global representation of each statistical feature vector. S-Net extracts the latent semantic representation of text input via a pre-trained RoBERTa model. G-Net aligns the information from the two sources and then adjusts the information flow.
Generative Adversarial Network
GAN-EDC (Duan et al., 2021) and Adanomaly (Qi et al., 2022) utilize technologies related to Generative Adversarial Network (GAN). GAN-EDC employs an LSTM-based encoder-decoder framework as the generator. Log keywords are taken as the input of the encoder, and the decoder outputs the generated log template. The discriminator utilizes CNN to identify the disparities between the generated log template and the real log template. Adanomaly is a log anomaly detection method based on Bidirectional GAN (BiGAN) and the stacking approach. It uses a CNN-based BiGAN model to extract features. Subsequently, it trains multiple classifiers by partitioning the extracted features into multiple subsets and integrates the classifiers using the stacking method.
Graph Neural Network
As data provenance graphs are complex datasets, typically represented as heterogeneous graphs and multiplex networks, PROV-GEM (Kapoor et al., 2021) employs a GCN-based approach. It is specially designed for attribute heterogeneous multiplex networks and is able to handle large networks and the heterogeneity of data provenance graphs.
@@ 5.3. Attack Investigation
Table 5. Overview of attack investigation approaches.
Approach Release Time Audit Log Applica- tion Log Use of Graph Base Model Starting Node Investigation Granularity
ProvDetector (Wang et al., 2020) 2020 - Doc2Vec - Path
ATLAS (Alsaheel et al., 2021) 2021 - LSTM Sequence
Liu et al. (Liu et al., 2022a) 2022 - Struc2Vec Provenance Graph
LogTracer (Niu et al., 2022) 2022 - DeepLog Path
ConLBS (Li et al., 2023b) 2023 - Transformer Provenance Graph
AirTag (Ding et al., 2023b) 2023 - BERT - Sequence
Karios (Cheng et al., 2024) 2023 GNN - Summary Graph
CAPTAIN (Ramaki et al., 2023) 2023 - - Provenance Graph
Except for identifying individual compromised nodes, IDS are supposed to unveil the full story of intrusions (a.k.a., attack scenarios). This process is referred to as attack investigation, which can be done by directly detecting attack scenarios (Wang et al., 2020), or analyzing the causal relations between compromised nodes (Alsaheel et al., 2021; Ding et al., 2023b). The attack scenarios are defined with scenario graphs as follows:
@@
(Attack Scenario). Attack scenario is a form of scenario graph, where its nodes of interest are compromised nodes.
In the past, attack investigation is conducted by forward analysis and backward analysis. Forward analysis discovers the influence that nodes of interest will cause and backward analysis traces back how nodes of interest are generated. Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns and predicting possible attack scenarios. Table 5 summarizes the overview of attack investigation approaches.
@@
ProvDetector (Wang et al., 2020) utilizes doc2vec to learn the embedding representation of paths in the provenance graph. This method enables the mapping of different components in the path (e.g., processes, files, and network connections) to a multidimensional numerical vector space. CAPTAIN (Ramaki et al., 2023) is a community-based APT analyzer in IT networks to detect Intrusion Kill Chain (IKC). CAPTAIN consists of a total of 12 different activities in two stages to detect possible IKCs of the APT attacks. CAPTAIN uses graph-based algorithms to model relationships among logged events. It detects event communities and summarizes them with centrality measures. By considering the nature of included events, it identifies malicious communities related to IKC stages. This ultimately leads to the reconstruction of the IKC of APT attacks.
Recurrent Neural Network
ATLAS (Alsaheel et al., 2021) is a framework that constructs end-to-end attack stories from readily available audit logs. It employs a novel combination of causal analysis, natural language processing, and machine learning techniques to build a sequence-based model. In ATLAS, LSTM enables the model to automatically learn to distinguish patterns in attack and non-attack sequences. This model establishes key patterns of attack and non-attack behaviors from causal graphs. LogTracer (Niu et al., 2022) is an efficient log-based anomaly tracing method. In attack path extraction, it weights edges of the initial provenance graph with an outlier algorithm, filters edges by threshold, and restructures the scenario graph. It finds the longest path as an attack path by assigning anomaly degrees to edges. It also uses an incremental graph processing strategy and the DeepLog model to process coarse-grained log information to assist the tracing process.
Transformer Architecture
AirTag (Ding et al., 2023b) employs unsupervised learning to train DL models directly from log texts rather than relying on labeled causal graphs. It uses the BERT model for unsupervised learning and retrains the BERT model based on unlabeled log data. ConLBS (Li et al., 2023b) combines a contrastive learning framework and multilayer Transformer network for behavior sequence classification. It constructs behavior sequences from audit logs, uses a lemmatization strategy to map semantics to the attack pattern layer, and can perform unsupervised representation learning on unlabeled sequences.
@@
Liu et al. (Liu et al., 2022a) propose an automated attack detection and investigation method via learning the context semantics of the provenance graph. The provenance graph analyzes temporal and causal dependencies of system events. The pre-trained Struc2Vec model is used to learn representations of unique nodes in a general set for node embedding. Kairos (Cheng et al., 2024) is a practical intrusion detection and investigation tool based on whole-system provenance. Kairos utilizes GNN to analyze system execution history and detect and reconstruct complex APTs. It employs a graph neural network-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event.
[…]
Figure 7. Taxonomy of attack investigation.
@@ 6.1. Dimensions of Datasets
To illustrate the quality of datasets, it is general to use the following dimensions:
@@ 6.2. Public Datasets
Loghub dataset (Zhu et al., 2023) is a large collection of system log datasets, providing 19 real-world log data from various software systems, including distributed systems, supercomputers, operating systems, mobile systems, server applications, and standalone software. The objective of Loghub is to fill the significant gap between intelligent automated log analysis techniques and successful deployments in the industry. For the usage scenarios of Loghub, about 35% are anomaly detection, 13% are log analysis, and 8% are security.
@@ 7.1.1. Poor Data Quality
Existing datasets for DL-IDS may contain errors, inaccuracies, or missing values. This leads to unreliable descriptions of system behaviors that may mislead DL-IDS. For example, in some cases of the DARPA TC dataset, the PROCESS object and its source fail to properly resolve conflicts, resulting in possible incorrect entries in the actor_id field. And the acuity_level value of the FLOW object is 0, while the value range for this field in other objects is from 1 to 5. Another example could be the LogChunks (Brandt et al., 2020) dataset. In this dataset, the content describing the failure reasons is possibly incomplete. This is because a chunk in LogChunks only contains a continuous substring of the log text and a failure reason may be described across multiple sections of the log. Moreover, LogChunks neglects the classification of failure reasons like test, compilation, and code inspection errors, which hinders further research from analyzing failure reasons.
@@
Meanwhile, high-quality ground-truth labels are hard to acquire, which is impeded by the contradiction between fine-grained manual labeling and automated label generation. On one hand, for unknown intrusions such as zero-day attacks, security analysts usually have no trouble recovering attack scenarios, while it is very labor-intensive to correspond each attack scenario to certain log entries. The DAPRA TC dataset (DARPA, 2022) is a typical example of this phenomenon. It only provides a ground truth report for attack scenarios, which does not correspond to any specific log entries. Although a few researchers (Wang et al., 2022a) provide the third-party ground-truth labels that are manually identified by themselves, we empirically find some ambiguities between their ground-truth labels and the official attack scenario report. These ambiguities have an obviously negative effect on DL-IDS, and to some extent, they may even cause the accumulation of errors. On the other hand, the development of automated labeling tools is in an awkward position. The automation requires prior knowledge of intrusions to design auto-labeling techniques such as instrumentation. In other words, the automation can only handle intrusions that it can understand and detect, making it relatively useless to develop such auto-labeling tools.
@@
In addition, there are no unified evaluation metrics for DL-IDS on these datasets (Cheng et al., 2024). For example, precision, recall, F1 score are usually exploited in most of work (Wang et al., 2020; Alsaheel et al., 2021; Rehman et al., 2024; Jia et al., 2024). Yet some papers (Ding et al., 2023b) propose to use True Positive Rate (TPR) and False Positive Rate (FPR) as evaluation metrics. All the comparison experiments between DL-IDS are thus hard to guarantee that the obtained performance is significant and practical. The potential of datasets will be weakened under this non-unified performance evaluation.
@@ 7.1.2. Insufficient Amount of Data
Although log data is generated very quickly (e.g., eBay generates 1.2 PB log data per day by 2018 (Samuel, 2018)), DL-IDS is still facing challenges in insufficient amounts of data. The reasons can be three-fold:
First, log data has an extremely large number of trivial events, which are proven to be not effective and usually removed by graph summarization (Zengy et al., 2022; Yang et al., 2023b). For example, data provenance provides fine-grained information about memory-related events, such as data-to-memory mapping and protection of certain memory addresses. These memory-related events basically do not involve attacks, and unfortunately, are always orthogonal to the existing DL-IDS. However, to ensure the completeness of data provenance and to capture very infrequent but inevitable memory attacks, these memory-related events are still recorded in benchmark datasets. Therefore, the usable part of each dataset is rather smalle for DL-IDS, which can be reflected by the high summarization ratio achieved by graph summarization approaches (e.g., 70% (Xu et al., 2016)).
@@
Table 7. Summarization of computating resources.
Type Approach Device Quantity
GPU DeepHunter (Wei et al., 2021a) Tesla P4 2
SHADEWATCHER (Zengy et al., 2022) Tesla V100 1
PROGRAPHER (Yang et al., 2023b) GeForce GTX TITAN X 1
FLASH (Rehman et al., 2024) GeForce RTX 2080 1
LogBD (Liu et al., 2023) GeForce RTX 2080 1
HDLNIDS (Qazi et al., 2023) GeForce GTX 2080Ti 1
NeuralLog (Le and Zhang, 2021) Tesla K40c 1
LogUAD (Wang et al., 2022b) Tesla V100 1
TCN-Log2Vec (Jiang et al., 2023) GeForce GTX 1660s 1
LogTracer (Niu et al., 2022) GeForce GTX 3080Ti 1
AIRTAG (Ding et al., 2023b) GeForce RTX 6000 1
MAGIC (Jia et al., 2024) GPU Unknown
CPU Only UNICORN (Han et al., 2020) vCPU 8
THREATRACE (Wang et al., 2022a) vCPU 16
Pagoda (Xie et al., 2018) Physical CPU 2
PROVDETECTOR (Wang et al., 2020) Physical CPU 1
SigML++ (Trivedi et al., 2023) Physical CPU 1
LogAnomaly (Meng et al., 2019) Physical CPU 1
GAN-EDC (Duan et al., 2021) Physical CPU 1
LogBASA (Liao et al., 2023) Physical CPU 1
KAIROS (Cheng et al., 2024) Physical CPU 1
CAPTAIN (Ramaki et al., 2023) Physical CPU 1
Paradise (Wu et al., 2022) Physical CPU Unknown

### v2

@@ 5.2.3. Network Traffic-based Detectors
Network traffic comes from communications between hosts across a computer network. It is ruled by network protocols such as Transmission Control Protocol (TCP) and the User Datagram Protocol (UDP) and can be utilized for intrusion detection. Basically, network traffic-based detectors are termed NIDS.
@@
Given the fact that network traffic is usually encrypted for secure communications, feature engineering-guided machine learning is widely applied in NIDS. Whisper (Fu et al., 2021) pays attention to both high accuracy and high throughput by utilizing frequency domain features. SigML++ (Trivedi et al., 2023) is an extension of SigML for supervised anomaly detection approach. SigML++ employs Fully Homomorphic Encryption and Artificial Neural Network (ANN) for detection, resulting in execution without decrypting the logs. OADSD (Zhang et al., 2023a) achieves task independently and has the ability of adapting to the environment over SD-WAN by using On-demand Evolving Isolation Forest. LtRFT (Tang et al., 2023) innovatively introduces Learning-To-Rank scheme for mitigating the low-rate DDoS attacks targeted at flow tables.
Graph and Sequence Neural Network
In network traffic, packets consist of various contents and their flows can be represented as graphs. As a result, both graph neural network and sequence neural network are adopted in NIDS. Kitsune (Mirsky et al., 2018) is a plug and play NIDS that is allowed to detect attacks efficiently on the local network without supervision. It alleviates the problem that network gateways and router devices simply do not have the memory or processing power. MT-FlowFormer (Zhao et al., 2022a) is a semi-supervised framework to mitigate the lack of a mechanism for modeling correlations between flows and the requirement of a large volume of manually labeled data. I2RNN (Song et al., 2022) is an incremental and interpretable RNN for encrypted traffic classification, which can be efficiently adapted for incremental traffic types. ERNN (Zhao et al., 2022b) represents error-resilient RNN, which is a robust and end-to-end RNN model specially designed against network-induced phenomena. Euler (King and Huang, 2023) accelerates the most memory-intensive part, message-passing stage within GNN, with several concurrently-executed replicated GNNs. pVoxel (Fu et al., 2023) is an unsupervised method that proposes to leverage point cloud analysis to reduce false positives for the previous NIDS such as Whisper and Kitsune without requiring any prior knowledge on the alarms. NetVigil (Hsieh et al., 2024) is specially designed for east-west traffic within data center networks. It utilizes E-GraphSage and contrastive learning techniques to strengthen its resilience. Exosphere (Fu et al., 2024) detects flooding attacks by analyzing packet length patterns, without investigating any information in encrypted packets. DFNet (Zhao et al., 2024a) is a DDoS prevention paradigm denoted by preference-driven and in-network enforced shaping. RFH-HELAD (Zhong et al., 2024) consists of a $K$ classification model based on a deep neural network and a $K+1$ classification combining GAN and Deep kNN for detecting anomalies in network traffic.
5.2.4. Hybrid Log-based Detectors
Based on the above discussions, a natural idea is to combine various types of log data for improving detection capability. OWAD (Han et al., 2023) is a general framework to detect, explain, and adapt to normality shifts in practice. OWAD is validated to be effective in various detection granularity, covering provenance graphs, application logs, and network packets. FG-CIBGC (Ni et al., 2025) mines syncretic semantics in multi-source logs including audit logs, application logs, and network traffic using LLM under in-context learning, which generates behavior graphs for comprehensive analysis.
@@ 5.3. Attack Investigation
Except for identifying individual intrusive nodes, IDS are supposed to detect the full story of intrusions (a.k.a., attack scenario graphs). This process is referred to as attack investigation, which can be done by directly detecting attack scenario graphs (Wang et al., 2020), or analyzing the causal relations between compromised nodes progressively to construct attack scenario graphs (Alsaheel et al., 2021; Ding et al., 2023b; Xu et al., 2022; Jiang et al., 2025). The attack scenario graphs are defined with scenario graphs as follows:
@@
(Attack Scenario Graph). Attack scenario graph is a scenario graph where its nodes of interest are compromised nodes.
In the past, attack investigation is conducted by forward analysis and backward analysis (Hossain et al., 2017). Forward analysis discovers the influence that nodes of interest will cause and backward analysis traces back how nodes of interest are generated. Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns of attack scenario graphs. Table 5 summarizes the overview of attack investigation approaches. Similar to Section 5.2, we exclude papers (Fang et al., 2022; Gao et al., 2018; Ma et al., 2017; Li et al., 2024a; Milajerdi et al., 2019b; Hassan et al., 2020b; Yang et al., 2020; Kwon et al., 2018; Alhanahnah et al., 2022; Ji et al., 2017; Zhu et al., 2023b; Hossain et al., 2017; Wang et al., 2023) slightly relevant to DL for conciseness.
Table 5. Overview of attack investigation approaches.
Taxonomy Approach Release Time Audit Log Applica- tion Log Base Model Starting Node Investigation Granularity
Traditional ProvDetector (Wang et al., 2020) 2020 doc2vec Path
Sequence ATLAS (Alsaheel et al., 2021) 2021 LSTM Graph
Neural LogTracer (Niu et al., 2022) 2022 DeepLog Path
Network ConLBS (Li et al., 2023a) 2023 Transformer Graph
AirTag (Ding et al., 2023b) 2023 BERT Graph
Graph Liu et al. (Liu et al., 2022a) 2022 struc2vec Graph
Neural Karios (Cheng et al., 2024) 2023 GNN Graph
Network TREC (Lv et al., 2024) 2024 GNN Graph
Orthrus (Jiang et al., 2025) 2025 UniMP Path
@@
Unlike detecting intrusive nodes, attack scenario graphs are complicated and thus are hard to handle by traditional learning methods. ProvDetector (Wang et al., 2020) utilizes doc2vec to learn the embedding representation of paths in the provenance graph. Then a density-based detection is deployed to detect abnormal causal paths in the provenance graph.
Sequence Neural Network
Log data is in the form of natural language text or is allowed to be transformed into sequences of events, which facilitates the introduction of sequence neural networks. ATLAS (Alsaheel et al., 2021) is a framework to construct end-to-end attack stories from readily available audit logs, which employs a novel combination of causal analysis and natural language processing. ATLAS exploits LSTM to automatically learn the pattern difference between attack and non-attack sequences. LogTracer (Niu et al., 2022) is an efficient anomaly tracing framework that combines data provenance and system log detection together. An outlier function with an abnormal decay rate is introduced to improve the accuracy. ConLBS (Li et al., 2023a) combines a contrastive learning framework and multilayer Transformer network for behavior sequence classification. AirTag (Ding et al., 2023b) employs unsupervised learning to train BERT directly from log texts rather than relying on provenance graphs. AirTag constructs attack scenario graphs by integrating the detected victim nodes.
@@
To capture causal relations within graphs, GNN is commonly adopted. Liu et al. (Liu et al., 2022a) propose an automated attack detection and investigation method via learning the context semantics of the provenance graph. The provenance graph analyzed by struc2vec captures temporal and causal dependencies of system events. Kairos (Cheng et al., 2024) is a practical intrusion detection and investigation tool based on whole-system provenance. Kairos utilizes GNN to analyze system execution history, so that detects and reconstructs complex APTs. It employs a GNN-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event. TREC (Lv et al., 2024) abstracts APT attack investigation problem as a tactics / techniques recognition problem. TREC trains its model in a few-shot learning manner by adopting a Siamese neural network. Orthurus (Jiang et al., 2025) identifies Quality of Attribution as the key factor contributing to whether or not the industry adopts IDS. It first detects malicious hosts using a GNN encoder and then reconstructs the attack path through dependency analysis.
@@ 6.1. Dimensions of Datasets
To illustrate the quality of DL-IDS datasets, it is general to use the following dimensions:
@@ 6.2. Public Datasets
LogChunks (Brandt et al., 2020) is an application log dataset for build log analysis, containing 797 annotated Travis CI build logs from 80 GitHub repositories and 29 programming languages. These logs are from mature and popular projects, collected through repository, build, and log sampling. Each log in the dataset has manually labeled text blocks of build failure reasons, search keywords, and structural categories, and cross-validated with the original developers with an accuracy of 94.4%.
Loghub dataset (Zhu et al., 2023a) is a large collection of system log datasets, providing 19 real-world log data from various software systems, including distributed systems, supercomputers, operating systems, mobile systems, server applications, and standalone software. The objective of Loghub is to fill the significant gap between intelligent automated log analysis techniques and successful deployments in the industry. For the usage scenarios of Loghub, about 35% are anomaly detection, 13% are log analysis, and 8% are security.
@@ 7.1.1. Poor Data Quality
Existing datasets for DL-IDS may contain errors, inaccuracies, or missing values. This leads to unreliable descriptions of system behaviors that may mislead DL-IDS. For example, in some cases of the DARPA TC dataset, the PROCESS object and its source fail to properly resolve conflicts, resulting in possible incorrect transformation. Besides, the acuity_level value of the FLOW object is 0, while the value range for this field in other objects is from 1 to 5. Another example could be the LogChunks (Brandt et al., 2020) dataset. In this dataset, the content describing the failure reasons is possibly incomplete. This is because a chunk in LogChunks only contains a continuous substring of the log text and a failure reason may be described across multiple sections of the log. Moreover, LogChunks neglects the classification of failure reasons like test, compilation, and code inspection errors, which hinders further research from analyzing failure reasons.
@@
Meanwhile, high-quality ground-truth labels are hard to acquire, which is impeded by the contradiction between fine-grained manual labeling and automated label generation. On one hand, for unknown intrusions such as zero-day attacks, it is very labor-intensive for security analysts to correspond each attack scenario to certain log entries, although coarse-grained attack scenarios may have been acquired. The DAPRA TC dataset (DARPA, 2022) is a typical example for this. It only provides a ground truth report for attack scenarios, which does not correspond to any specific log entries. Although a few researchers (Wang et al., 2022) provide the third-party ground-truth labels that are manually identified by themselves, we empirically find some ambiguities between their ground-truth labels and the official attack scenario report. These ambiguities have an obviously negative effect on DL-IDS, and to some extent, they may even cause the accumulation of errors. On the other hand, the development of automated labeling tools is in an awkward position. The log data is generated based on its given prior knowledge of intrusions (Cheng et al., 2025), whereas the challenge of DL-IDS is to detect zero-day intrusions. This tends the development of such automated tools to be somewhat pointless.
@@
In addition, there are no unified and effective evaluation metrics for DL-IDS (Cheng et al., 2024), which further weakens the potential of datasets. For example, precision, recall, F1 score are usually exploited in most studies (Wang et al., 2020; Alsaheel et al., 2021; Rehman et al., 2024; Jia et al., 2024), while some papers (Ding et al., 2023b) propose to use True Positive Rate (TPR) and False Positive Rate (FPR) as evaluation metrics. This makes the comparison experiments usually unfair and hard to tell if the validation is convincing. We also note that in many cases where the percentage of negatives (or malicious log entries) is low, sacrificing FPR can always significantly increase TPR. For example, sacrificing 1,000 false positives for one true positive might only increase FPR by 0.05%, but would increase TPR by 5%.
@@ 7.1.2. Insufficient Amount of Data
Although log data is generated very quickly (e.g., eBay generates 1.2 PB log data per day by 2018 (Samuel, 2018)), DL-IDS is still facing challenges in insufficient amounts of data. Discounting the above data quality issues such as inaccuracies, the reasons are three-fold:
First, log data has an extremely large number of trivial events, which are proven ineffective and usually removed by graph summarization (Zengy et al., 2022; Yang et al., 2023). For example, data provenance provides fine-grained information about memory-related events, such as data-to-memory mapping and protection of certain memory addresses. These memory-related events basically do not involve attacks, and unfortunately, are always orthogonal to the existing DL-IDS. However, to ensure the completeness requirement of data provenance and to capture very infrequent but inevitable memory attacks, these memory-related events are still recorded in benchmark datasets. As a result, the usable part of each dataset is rather small for DL-IDS, which can be reflected by the high summarization ratio achieved by graph summarization approaches (e.g., 70% (Xu et al., 2016)).